hello, how do i solve this please:
(1/a + 1/b)x=1 for x
I would like to learn the steps as well.
Thanks.

Found 2 solutions by richard1234, richwmiller:
Answer by richard1234(7193)

Multiply both sides by (1/a + 1/b):



The denominator can be simplified by using the GCD:


Answer by richwmiller(17219)

you want to leave just x on one side.
Divide both sides by (1/a + 1/b)
x=1/(1/a + 1/b)
or
(1/a + 1/b)x=1
x/a+x/b=1
get common denominator
xb/ab+xa/ab=1
(xb+xa)/ab=1
x(a+b)/ab=1
1/x=(a+b)/ab
x=ab/(a+b)




(ab)/(a+b) = x